"Native," or elemental copper can be found in nature, but most copper is mined as oxide or sulfide minerals. Chalcopyrite (CuFeS2) is one copper mineral that can be converted to elemental copper in a series of chemical steps. Reacting chalcopyrite with oxygen at high temperature produces a mixture of copper sulfide and iron oxide. The iron oxide is separated from CuS by reaction with sand (SiO2). CuS is converted to Cu2S in the process and the Cu2S is burned in air to produce Cu and SO2:

A)An average copper penny minted in the 1960s contained about 3.000 g of copper. How much chalcopyrite had be mined to produce 100 pennies?

b)How much chalcopyrite had to be mined to produce 100 pennies if reaction 1 had a percent yield of 90.00 % and all other reaction steps had yield of 100%?

c)How much chalcopyrite had be mined to produce 100 pennies if each reaction involving copper proceeded in 90.00 % yield?

"Native," or elemental copper can be found in nature, but most copper is mined as oxide or sulfide minerals. Chalcopyrite (CuFeS2) is one copper mineral that can be converted to elemental copper in a series of chemical steps. Reacting chalcopyrite with oxygen at high temperature produces a mixture of copper sulfide and iron oxide. The iron oxide is separated from CuS by reaction with sand (SiO2). CuS is converted to Cu2S in the process and the Cu2S is burned in air to produce Cu and SO2:

A)An average copper penny minted in the 1960s contained about 3.000 g of copper. How much chalcopyrite had be mined to produce 100 pennies?

b)How much chalcopyrite had to be mined to produce 100 pennies if reaction 1 had a percent yield of 90.00 % and all other reaction steps had yield of 100%?

c)How much chalcopyrite had be mined to produce 100 pennies if each reaction involving copper proceeded in 90.00 % yield?
Answer #1

(c) If all the steps have 90% yield, then over all yield is : (0.9 * 0.9 *0.9 *0.9) * 100% = 65.61 %

mass of chalopyrite to be minted = 865 * 100/ 65.61 g = 1318 g

do rate it if you are satisfied with answer